OpenStudy (heart_offire):

Simplify the expression. x2 - 9 ________ x2 + 2x - 15 A) x + 3 ______ x + 5 B) 3 _____ 2x + 5 C) x - 1 _____ x - 6 D) 9 _____ 15 - 2x @study312

OpenStudy (study312):

One sec

OpenStudy (heart_offire):

im confused on this because there are 2 numbers on top and three on the bottom

OpenStudy (study312):

just add both

OpenStudy (study312):

2 - 9 = -7 2 + 2 = 4, 4 - 15 = -11 -7 ÷ -11 = 0.64

OpenStudy (study312):

so I am actually kinda lost because the top would be -7 unless it wants you to leave the 2x and x2 alone

OpenStudy (heart_offire):

maybe....i dont know because sometimes these problems are like that.

OpenStudy (study312):

is x2 \[x ^{2}\]?

OpenStudy (study312):

@Ms-Brains do you think you could solve x2 - 9 and 2x + x2 - 15?

MsBrains (ms-brains):

http://www.cymath.com/answer.php?q=factor%28x%5E2-9%29 Use this this site guys!

OpenStudy (study312):

thanks Ms-Brains

OpenStudy (heart_offire):

@study yes....sorry...yes it is an exponent

OpenStudy (heart_offire):

i forgot to put these x^2

MsBrains (ms-brains):

No probs cx

OpenStudy (heart_offire):

i think the answer is A

OpenStudy (study312):

Yup its A